OpenStudy (girlgoyleh.):

The system of equations is inconsistent. What are the missing values? 5x+___y=2 ___x+3y=8

OpenStudy (danjs):

inconsistent means there will not be a solution to the system of two lines. In other words, the two lines will not intersect, and are parallel.

OpenStudy (girlgoyleh.):

how would i find the answers to the blank spaces then?

OpenStudy (danjs):

An easy way to determine if two lines are parallel from the given form above... Ax + By = E Cx + Dy = F If the constants are a multiple of one another than the lines are parallel. If you can multiply, Ax+By by some number , and get Cx+Dy, the lines will be parallel.

OpenStudy (danjs):

So if you let the first equation be... 5x + 1y = 2 you can see that the second equation will be 3 times the first, The first will have 1y, and the second has 3y. So you have to multiply the 5x by 3

OpenStudy (danjs):

3*(5x + 1y) ---> 15x + 3y Here is a good reference for these types of probs... http://www.algebra.com/algebra/homework/coordinate/Types-of-systems-inconsistent-dependent-independent.lesson

OpenStudy (girlgoyleh.):

thanks for explaining that :)

OpenStudy (danjs):

A step further, if you multiply the first equation by a constant number, and not only the values on the X and Y become the same, but also the value on the Right of the = sign becomes the same, then Both are the same line. For example 4X + 5Y = 5 8X + 10Y =10 The second equation is just 2 times the first equation. Both are the same line, they overlap. They have Infinitely many solutions, and are dependent.

OpenStudy (girlgoyleh.):

yeah, it was actually 3 and 5

OpenStudy (danjs):

that is possible too. As long as the coefficients on X and Y in the second line are a multiple of the coefficients of X and Y in the second line. Using 3 and 5, means that they used a multiple of 1.